OpenStudy (samsan9):

what is the vertex form of y=(x+2)(x-3)?

OpenStudy (anonymous):

y=(x+2)(x-3) y=x^2 - x -6 try completeing squares first

OpenStudy (anonymous):

First expand: \[(x+2)(x-3)=x^2-x-6\] Complete the square to write the parabola in vertex form: \[x^2-x+\frac{1}{4}-\frac{25}{4}=\cdots\]

OpenStudy (samsan9):

so you = it to zero and add 6 to both sides

OpenStudy (anonymous):

no see what SithsAndGiggles has done for completing the squares y =x^2 - x -6 y = x^2 -2*(1/2) x +1/4 -1/4 -6 y=(x-1/2)^2 -25/4 or y+25/4 = (x-1/2)^2 now i hoep you may be able to sate the vertex

OpenStudy (samsan9):

so the vertex of the parabola right?

OpenStudy (anonymous):

1/2 ,-25/4

OpenStudy (samsan9):

so now what ?

OpenStudy (triciaal):

The vertex form of a parabola's equation is generally expressed as : y= a(x-h)2+k where (h, k) is the vertex

OpenStudy (samsan9):

so do i plug in the 1/2 and -25/4 into the equation?

OpenStudy (triciaal):

yes y = (x-1/2) -25/4

OpenStudy (triciaal):

another approach after you expand and you have y = x^2-x - 6 a = 1 b = -1 and c = -6 the h value for the vertex = -b/2a = -(-1)/2*1 = 1/2 the k value is f(h) = (1/2)^2 -1/2 - 6 = 1/4 -1/2 -6 = -25/4

OpenStudy (samsan9):

so how do i put it into vertex form?

OpenStudy (triciaal):

repost yes y = (x-1/2)^2 -25/4